OpenStudy (anonymous):

The Positive Square Root Of 11 Is Between

OpenStudy (anonymous):

What's 3 squared?

hartnn (hartnn):

think about 2 perfect squares , one less than 11 and other greater than 11

OpenStudy (anonymous):

So Its 4 And 5 Right

hartnn (hartnn):

nopes, 3^2 = 9 and 4^2 = 16 so...?

OpenStudy (anonymous):

Ohhh Okay Hold On I Believe Its 5 And 6 But Wait

OpenStudy (anonymous):

How close you have to get?

OpenStudy (anonymous):

Yes Its 5 AND 6

hartnn (hartnn):

5^2 =25 6^2 =36 11 is not between them....

OpenStudy (anonymous):

Okay Now I Get It Hold On Wait A Minute

OpenStudy (anonymous):

Its 3 And 5

hartnn (hartnn):

3^2 = 9 and 4^2 = 16 11 is between them.....

hartnn (hartnn):

so square root of 11 will be between \(\sqrt{3^2}and \sqrt{4^2}\)

hartnn (hartnn):

what is \(\sqrt{a^2}=?\)

OpenStudy (anonymous):

Thankks I Got It

hartnn (hartnn):

did u get it as between 3 and 4 ?? is that an option ?

OpenStudy (anonymous):

Yes